# Perspective on Physical Interpretations of Rényi Entropy in Statistical Mechanics

### v1

@@ I Introduction
In quantum-many-body systems for instance the Rényi entropy gives a measure of quantum entanglement. It can be measured in experiments [9, 10] and in numerical simulations that are less computationally demanding [11, 12, 13, 14, 15]. The use of the replica trick and conformal field theory allow moreover to compute analytically Rényi entanglement entropy for some systems, and to obtain the (Shannon or Von Neumann) entanglement entropy by analytical continuation [16]. In dynamical systems, it is common practice to use the Rényi entropy as a characterization of chaos [17, 18]. The Rényi entropy is also related to a theoretical formalism in non-equilibrium dynamics. It is a generalization of the Kolmogorov-Sinai entropy, which corresponds to the cumulant generating function associated with a dynamical partition function involving counting trajectories in phase space [19, 20]. Since the Rényi entropy is tightly related to the replica trick [21, 22], it has relationships with the physics of disordered systems. Reference [23] connects the configurational entropy characterizing a thermodynamic glass transition [24, 25] and real-space structure through the Rényi entropy. For other applications in physics, see Ref. [26] and references therein.
@@ II Derivation of Rényi entropy
As shown in the introduction, the Rényi entropy with various values of $q$ has been widely used in physics. One of the main goals of this paper is to provide concise interpretations of the Rényi entropy with index $q$ in physical systems.
III General properties of Rényi entropies
@@
the collision entropy, that can be interpreted as follows. Suppose that we perform two independent trials generated by the same probability distribution (denoted as $p_{i}$). The probability that the two trials give rise to the same specific event, say, an event $i$, is $(p_{i})^{2}$. Thus $\sum_{i}(p_{i})^{2}$ corresponds to the probability of observing the same event irrespective of the type of event. $S_{2}^{\rm R\acute{e}nyi}$ in Eq. (8) is (the minus of) the logarithm of such probability. Alternatively, one can think about a joint system composed of two identical but independent replicas following the same probability distribution. $S_{2}^{\rm R\acute{e}nyi}$ corresponds to the logarithmic of the probability that the two replicas give rise to the same event. The above interpretations can be easily generalized to higher (integer) values of $q$: $S_{q}^{\rm R\acute{e}nyi}$ in Eq. (4) corresponds to the logarithm of the probability of observing the same event for all of $q$ replicas (which are identical yet independent) . The notion of replicas arises therefore naturally in the Rényi entropy, and makes it easier to measure in simulations and experiments than the Shannon entropy. This is one of the main reasons why it has been used in various contexts, in particular dynamical systems [17, 18] and quantum many-body systems [9, 10]. In the next sections we discuss further this connection with replicas.
@@
where $\tilde{\bf p}^{q}=(\tilde{p}_{1}^{q},\tilde{p}_{2}^{q},\cdots,\tilde{p}_{\Omega}^{q})$ is a normalized probability distribution whose component is given by $\tilde{p}_{i}^{q}=(p_{i})^{q}/\sum_{j}(p_{j})^{q}$ (and hence $\sum_{i}\tilde{p}_{i}^{q}=1$), and $D_{\rm KL}$ is the (standard) Kullback-Leibler divergence [2] which is non-negative. Therefore, $S_{q}^{\rm R\acute{e}nyi}$ is a non-increasing function of $q$, satisfying $S_{0}^{\rm R\acute{e}nyi}\geq S_{q}^{\rm R\acute{e}nyi}\geq S_{\infty}^{\rm R\acute{e}nyi}$, and $S_{q}^{\rm R\acute{e}nyi}$ is bounded by $S_{\infty}^{\rm R\acute{e}nyi}$ from below. Moreover, when $q>1$ one can also derive an upper bound of $S_{q}^{\rm R\acute{e}nyi}$ by $S_{\infty}^{\rm R\acute{e}nyi}$. Since $\sum_{i}(p_{i})^{q}\geq\max_{i}\{(p_{i})^{q}\}=(\max_{i}\{p_{i}\})^{q}$, and using Eqs. (4) and (9), gives $(q-1)S_{q}^{\rm R\acute{e}nyi}\leq qS_{\infty}^{\rm R\acute{e}nyi}$. Thus, when $q>1$,
@@
IV Rényi entropy for the Gibbs-Boltzmann distribution
@@
We note that the coupling between replicas is very strong, such that all different replicas condense in the same configuration. As an alternative to Eq. (21), one can also consider an external potential with a master-slave architecture, e.g., $i_{1}$ is the master that interacts with all slaves, $i_{2},i_{3},\cdots,i_{q}$, and there is no interaction between slaves. One can also consider fully connected interactions, i.e., all $q$ replicas interact with each other, akin to the replica method in disordered systems [22].
@@
where $I(e)$ is the rate function, and $N$ is the number of constituent elements (e.g., particles, spins). When $N\to\infty$, the mean value of the energy $\hat{e}$ converges to $\langle\hat{E}\rangle/N$, with probability one (law of large numbers), which corresponds to the minimization of $I(e)$ at $I(e)=0$. The variance of $\hat{e}$ would converge to zero with $N^{-1}$ (central limit theorem). The rate function $I(e)$ contains further information about large deviations [31]. Below we will show that the Rényi entropy contains the same information as $I(e)$ [44].
@@
V Discussion and Conclusions
The Rényi entropy is a one-parameter generalization of the Shannon entropy obtained by considering a generalized mean of the information content that preserves additivity (extensivity) for independent events, a fundamental aspect in information theory and a natural characteristic feature in many physical systems. Besides defining the Rényi entropy, statistical mechanics using this generalized expectation value may allow us to investigate large deviation, and possibly replica theory in a formally elegant manner [45].
The Rényi entropy includes various entropies used in physics as specific limits of $q$. In general, varying $q$ from the Shannon entropy limit $q\to 1$ corresponds to biasing or unbiasing the probability distribution. In particular, the Rényi entropy with a large value of $q$ is related to rare or atypical lower energy configurations. We reviewed physical interpretations of the Rényi entropy based on a statistical mechanical setting. The interpretations were provided by using free energy differences, the notion of replicas, work, and large deviation.
Estimating the Shannon entropy from the Rényi entropy is a practical important issue. Measurement of the Rényi entropy for particular integer values (e.g., $q=2$ and $3$) is easier than measuring the Shannon entropy. Developing reliable methods to estimate $q\to 1$ from $q=2,3,4,...$ is an important subject [33, 46].

### v2

@@ 1 Introduction
In quantum-many-body systems for instance the Rényi entropy gives a measure of quantum entanglement. It can be measured in experiments [9, 10] and in numerical simulations that are less computationally demanding [11, 12, 13, 14, 15]. The use of the replica trick and conformal field theory allow moreover to compute analytically Rényi entanglement entropy for some systems, and to obtain the (Shannon or Von Neumann) entanglement entropy by analytical continuation [16]. In dynamical systems, it is common practice to use the Rényi entropy as a characterization of chaos [17, 18]. The Rényi entropy is also related to a theoretical formalism in non-equilibrium dynamics. It is a generalization of the Kolmogorov-Sinai entropy, which corresponds to the cumulant generating function associated with a dynamical partition function involving counting trajectories in phase space [19, 20]. Since the Rényi entropy is tightly related to the replica trick [21, 22], it has relationships with the physics of disordered systems. Reference [23] connects the configurational entropy characterizing a thermodynamic glass transition [24, 25] and real-space structure through the Rényi entropy [26]. For other applications in physics, see Ref. [27] and references therein.
@@ 2 Derivation of Rényi entropy
As shown in the introduction, the Rényi entropy with various values of $q$ has been widely used in physics, such as a measure of entanglement in quantum systems (e.g., $q=2$  [9, 10]), in non-equilibrium statistical mechanics ($q=1/2$ and $q\to\infty$[33, 34], multi-fractal analysis [35], characterization of chaos ($q=2$[17], and as an estimate of the Shannon entropy (from $q=2,3,...$[36]. One of the main goals of this paper is to provide concise interpretations of the Rényi entropy with index $q$ in physical systems.
3 General properties of Rényi entropies
@@
the collision entropy, that can be interpreted as follows. Suppose that we perform two independent trials generated by the same probability distribution (denoted as $p_{i}$). The probability that the two trials give rise to the same specific event, say, an event $i$, is $(p_{i})^{2}$. Thus $\sum_{i}(p_{i})^{2}$ corresponds to the probability of observing the same event irrespective of the type of event. $S_{2}^{\rm R\acute{e}nyi}$ in Eq. (8) is (the minus of) the logarithm of such probability. Alternatively, one can think about a joint system composed of two identical but independent replicas following the same probability distribution. $S_{2}^{\rm R\acute{e}nyi}$ corresponds to the logarithmic of the probability that the two replicas give rise to the same event. The above interpretations can be easily generalized to higher (integer) values of $q$: $S_{q}^{\rm R\acute{e}nyi}$ in Eq. (4) corresponds to the logarithm of the probability of observing the same event for all of $q$ replicas (which are identical yet independent). The notion of replicas arises therefore naturally in the Rényi entropy, and makes it easier to measure in simulations and experiments than the Shannon entropy. Indeed $\sum_{i}(p_{i})^{q}$ in the Rényi entropy can be evaluated by making $q$ copies of the system, while the logarithmic form, $\sum_{i}p_{i}\log p_{i}$ in the Shannon entropy is not straightforward to measure. This is one of the main reasons why it has been used in various contexts, in particular dynamical systems [17, 18] and quantum many-body systems [9, 10]. In the next sections we discuss further this connection with replicas (see also Ref. [23] for a detailed discussion).
@@
where $\tilde{\bf p}^{q}=(\tilde{p}_{1}^{q},\tilde{p}_{2}^{q},\cdots,\tilde{p}_{\Omega}^{q})$ is a normalized probability distribution whose component is given by $\tilde{p}_{i}^{q}=(p_{i})^{q}/\sum_{j}(p_{j})^{q}$ (and hence $\sum_{i}\tilde{p}_{i}^{q}=1$), and $D_{\rm KL}$ is the (standard) Kullback-Leibler divergence [2] which is non-negative. Therefore, $S_{q}^{\rm R\acute{e}nyi}$ is a non-increasing function of $q$, satisfying $S_{0}^{\rm R\acute{e}nyi}\geq S_{q}^{\rm R\acute{e}nyi}\geq S_{\infty}^{\rm R\acute{e}nyi}$, and $S_{q}^{\rm R\acute{e}nyi}$ is bounded by $S_{\infty}^{\rm R\acute{e}nyi}$ from below. Moreover, when $q>1$ one can also derive an upper bound of $S_{q}^{\rm R\acute{e}nyi}$ by $S_{\infty}^{\rm R\acute{e}nyi}$. Since $\sum_{i}(p_{i})^{q}\geq\max_{i}\{(p_{i})^{q}\}=(\max_{i}\{p_{i}\})^{q}$, and using Eqs. (4) and (9), we get $(q-1)S_{q}^{\rm R\acute{e}nyi}\leq qS_{\infty}^{\rm R\acute{e}nyi}$. Thus, when $q>1$,
@@
4 Rényi entropy for the Gibbs-Boltzmann distribution
@@
We note that the coupling between replicas is very strong, such that all different replicas condense in the same configuration   [^5] As an alternative to Eq. (21), one can also consider an external potential with a master-slave architecture, e.g., $i_{1}$ is the master that interacts with all slaves, $i_{2},i_{3},\cdots,i_{q}$, and there is no interaction between slaves. One can also consider fully connected interactions, i.e., all $q$ replicas interact with each other, akin to the replica method in disordered systems [22]..
@@
where $I(e)$ is the rate function, and $N$ is the number of constituent elements (e.g., particles, spins). When $N\to\infty$, the mean value of the energy $\hat{e}$ converges to $\langle\hat{E}\rangle/N$, with probability one (law of large numbers), which corresponds to the minimization of $I(e)$ at $I(e)=0$. The variance of $\hat{e}$ would converge to zero with $N^{-1}$ (central limit theorem). The rate function $I(e)$ contains further information about large deviations [37]. Following Refs. [23, 49], we will show that the Rényi entropy contains the same information as $I(e)$.
@@
5 Discussion and Conclusions
The Rényi entropy is a one-parameter generalization of the Shannon entropy obtained by considering a generalized mean of the information content that preserves additivity (extensivity) for independent events, a fundamental aspect in information theory and a natural characteristic feature in many physical systems. Besides defining the Rényi entropy, statistical mechanics using this generalized expectation value allow to investigate large deviation, and possibly replica theory in a formally elegant manner [50].
@@ 5 Discussion and Conclusions
We also restricted to systems with discrete events, such as spins. Extending to continuous systems, such as classical particle systems, is seemingly straightforward. Yet it involves additional diverging terms (e.g., the logarithm of the bin size for discretization) in the same way as the Shannon entropy [56], which requires careful analysis for the system under investigation.